OpenStudy (anonymous):

simplify (4x^3y^5)^2

OpenStudy (callisto):

\[(4x^3y^5)^2\]\[=(4)^2(x^3)^2(y^5)^2\]\[=...\]

OpenStudy (anonymous):

Use power of a power property to finish the job.

OpenStudy (anonymous):

\[(a^n)^m=a^{nm}\]

OpenStudy (anonymous):

answer would be 16x^6y^10?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

awesome ok: Find each real number root: 1) \[\sqrt{-16}\] 2)\[\sqrt[3]{27}\]

OpenStudy (callisto):

New post for new question :\

OpenStudy (anonymous):

1 is "4i" 2, c'mon which number multiplying itself three times gives 27 (a*a*a=27)

OpenStudy (callisto):

4i is not a real number though.

OpenStudy (anonymous):

i is imaginary